# Recognition of Unit Segment and Polyline Graphs is $\exists\mathbb{R}$-Complete

### v2

@@ 2.1 Cell Lemma
We start by introducing the notion of a connector. Intuitively, it is just a bunch of vertices that separates an induced cycle from the (connected) rest of the graph.
@@
Let $G=(V,E)$ be a graph, and $C\subseteq V$ be a set of vertices forming an induced cycle, such that $G[V\setminus C]$ is connected. Let $D\subseteq V\setminus C$ be a set of vertices with the following properties.
  • The neighborhood of $C$ is $D$.
@@
Then we call the set $D$ connectors of $C$.
This definition is illustrated in Figure 6.
@@
Let $C$ be an induced cycle of $G=(V,E)$ with connectors $D$ and $r(V)$ be a geometric representation by Jordan arcs, then either $r(V\setminus(C\cup D))$ lies completely inside the interior curve of $r(C)$ or completely outside the exterior curve of $r(C)$.
@@
This follows from the fact that $G[V\setminus(C\cup D)]$ is connected and is not adjacent to the induced cycle $C$. Thus it is also not intersecting the interior or exterior boundary curve. By the Jordan Curve Theorem, the interior (exterior) boundary curve splits the plane in two and thus $G[V\setminus(C\cup D)]$ can only be in one of the two components. ∎
Figure 6: Left: The yellow dots form an induced cycle. The green vertices are connectors of this cycle and the remaining white squares form the rest of the graph. Right: The yellow dots form an induced cycle. The green dots are intersectors of this cycle. The cyclic order $o_{G}$ as used in Lemma 10 is $d_{2},d_{2},d_{1},d_{4},d_{3},d_{1}$, when starting at the top left of the cycle and going clockwise.
@@ 2.2 Order Lemma
The aim of this section is to show that certain vertices neighboring an induced cycle also need to intersect this cycle in the same order in any geometric representation. We start with a definition of the setup.
@@
Let $G=(V,E)$ be a graph, and $C\subset V$ a set of vertices forming an induced cycle. Let $D\subset V\setminus C$ be a set of vertices with the following properties.
@@
  • If $d\in D$ has two neighbors in $C$, these are non-adjacent.
@@
Then we call the set $D$ intersectors of $C$.
This definition is illustrated in Figure 6. We now define two cyclic orders of intersections of intersectors with the cycle. The first order lives in the realm of the graph, while the second one is concerned with a concrete realization. The goal of this section is to prove that these orders are the same.
@@
Let $G$ be a graph with $C$ an induced cycle on at least four vertices, and let $D$ be intersectors of $C$. When we travel along the cycle $C$ for one full rotation, we can write down the pairs $(c,d)\in C\times D$, such that $\{c,d\}\in E(G)$. This string of pairs defines the graph order of the intersectors up to a cyclic shift and reflection.
@@
Let $G$ be a graph with $C$ an induced cycle on at least four vertices, and let $D$ be intersectors of $C$. Let $r$ be a realization of $G$ by Jordan arcs and let $b$ be the interior boundary curve of $r(C)$. When we travel along $b$ for one full rotation, we can write down the pairs $(c,d)\in C\times D$ such that $r(d)$ intersects $b$ where $b$ is tracing $r(c)$. From consecutive copies of the same pair we only keep one. This string of pairs defines the geometric order of the intersectors up to a cyclic shift and reflection.
The graph order is illustrated in Figure 6, while the geometric order is illustrated in Figure 7.
@@
Let $r$ be a realization of a graph $G$ with an induced cycle $C$ and intersectors $D$. If for every intersection of some $r(d)$ for $d\in D$ with some $r(c)$ for $c\in C$ we have that $r(d)$ also intersects the interior boundary curve $b$ close to that intersection, then the geometric and graph order of the intersectors are the same.
Figure 7: A realization of an induced cycle $C$ (yellow Jordan arcs) and intersectors of $C$ (green Jordan arcs). The cyclic order $o_{r}$ along the interior boundary curve as used in Lemma 10 is here $(c_{5},d_{2}),(c_{7},d_{1}),(c_{2},d_{1})$, when starting at the bottom right and going clockwise (see black arrow).
@@
We first claim that every pair $(c,d)$ occurs in both orders exactly once if $\{c,d\}\in E(G)$, and zero times otherwise: In the graph order this holds by definition. Furthermore, the assumption of this lemma guarantees that any pair $(c,d)$ in the graph order of intersectors also occurs in the geometric order of intersectors. We thus only need to show that in the geometric order no pair occurs multiple times. A pair could only occur more than once if there is another pair $(c^{\prime},d^{\prime})$ between these occurrences. However, since $c$ is only crossed by $d$, and no neighbor of $c$ in $C$ is crossed by any intersector, Lemma 4 guarantees that no such pair $(c^{\prime},d^{\prime})$ can occur between the two occurrences of $(c,d)$.
Next, we show that the pairs are also ordered the same way. Since for any two pairs $(c,d),(c^{\prime},d^{\prime})$ in the two vertices $c,c^{\prime}$ are distinct and non-adjacent, Lemma 4 guarantees that the geometric order of the intersectors must respect the ordering of the $c_{i}$ along the cycle. Similarly, the graph order of intersectors must respect this ordering by definition. Thus, the two orders are the same. ∎
Note that the definition of the geometric order and Lemma 10 also work for the exterior boundary curve instead of the interior one.
@@ 3 $\exists\mathbb{R}$-Membership
In this section, we show the $\exists\mathbb{R}$-membership parts of Theorems 1 and 2.
@@
We describe the naive approach only for unit segments, however a similar technique also easily works for polylines. Let $G=(V,E)$ be a graph. For each vertex $v\in V$, we use four variables $\left\llbracket v1\right\rrbracket,\left\llbracket v2\right\rrbracket,\left\llbracket v3\right\rrbracket,\left\llbracket v4\right\rrbracket$ meant to describe the coordinates of the endpoints of a unit segment realizing $v$. We can construct ETR-formulas unit and intersection that test whether a segment has unit length and whether two segments are intersecting, respectively. The formula $\varphi$ consists of the three parts:
@@
The unit formula can be constructed using the formula for the Euclidean norm. To construct the intersection formula it is possible to use the orientation test: The orientation test formula checks whether a given ordered triple of points is oriented clockwise, or counter-clockwise. Using multiple orientation tests on the endpoints of two segments one can determine whether the segments intersect. The orientation test itself can be constructed using a standard determinant test. This finishes the description of the formula $\varphi$ and establishes $\exists\mathbb{R}$-membership.
Although all of these formulas are straightforward to describe, things get a bit lengthy (especially in the case of polylines) and we do hide some details about the precise polynomials. We therefore also wish to describe the second approach using real witnesses and verification algorithms. To use this approach we need to first introduce a different characterization of the complexity class $\exists\mathbb{R}$. Namely, an algorithmic problem is in $\exists\mathbb{R}$ if and only if we can provide a real verification algorithm [19]. A real verification algorithm $A$ for a problem $P$, takes as input an instance $I$ of $P$ and a polynomial-size real-valued witness $w$. $A$ must have the following properties: In case that $I$ is a yes-instance, there exists a $w$ such that $A(I,w)$ returns yes. In case that $I$ is a no-instance $A(I,w)$ will return no, for all possible $w$. Note that this is reminiscent of the definition of NP using a verifier algorithm. There are two subtle differences: The first one is that $w$ is allowed to contain real numbers and discrete values, not only bits. The second difference is that $A$ runs on the real RAM instead of a Turing machine. This is required since a Turing machine is not capable of dealing with real numbers. It is important to note that $I$ itself does not contain any real numbers. We refer to Erickson, Hoog, and Miltzow [19] for a detailed definition of the real RAM.
Given this alternative characterization of $\exists\mathbb{R}$, it is now very easy to establish $\exists\mathbb{R}$-membership of Unit Recognition and $k$-PolyLine Recognition: We merely need to describe the witness and the verification algorithm. The witness is a description of the coordinates of the unit segments (or polylines, respectively) realizing the given graph. The verification algorithm merely checks that the correct realizations of vertices intersect one another, and in the case of unit segments, also checks that all segments have the correct length. Note that we sweep many details of the algorithm under the carpet. However, algorithms are much more versatile than formulas and it is a well-established fact that algorithms are capable of all types of elementary operations needed to perform this verification.
@@ 4 $\exists\mathbb{R}$-Hardness
In this section we prove $\exists\mathbb{R}$-hardness, first for unit segments, then for polylines. The reduction for polylines builds upon the reduction for unit segments and we will only highlight the differences.
@@ 4.1 $\exists\mathbb{R}$-Hardness for Unit Segments
We show $\exists\mathbb{R}$-hardness of Unit Recognition by a reduction from Stretchability.
@@
We present the reduction in three steps. First, we show how we construct a graph $G$ from a pseudoline arrangement $\mathcal{A}$. Then we show completeness, i.e., we show that if $\mathcal{A}$ is stretchable then $G$ can be represented using unit segments. At last, we will show soundness, i.e., we show that if $G$ can be represented using unit segments then $\mathcal{A}$ is stretchable. For this part we will use the lemmas from Section 2.
@@
Given a pseudoline arrangement $\mathcal{A}$ of $n$ pseudolines, we construct a graph $G$ by enhancing the arrangement $\mathcal{A}$ with additional Jordan arcs. Then we define $G$ to be the intersection graph of the pseudolines (which are also Jordan arcs) and all our added arcs. See Figure 8 for an illustration.
First, we add so called probes to our pseudolines. A probe of depth $k$ of pseudoline $\ell$ is a Jordan arc which starts on the left vertical line, and follows $\ell$ closely through the arrangement until it has intersected $k$ other pseudolines (and all probes of that other pseudoline which reach that intersection). For each pseudoline $\ell$, we add $2(n-1)$ probes: Above and below $\ell$ we add one probe each for each depth $11\leq k\leq n-$ 1. This gives a total of $p:=2n(n-1)$ probes in the graph $G$. The probes are sorted according to their depth, with the probes of smallest depth situated closest to $\ell$. Note that so far we have added $n+p$ arcs. We now create connectors for each probe and pseudoline. A connector is a short Jordan arc added to the left and/or right end of another arc. For probes, we only add connectors at the left end, while pseudolines get connectors at both ends. Note that we thus add $d:=2n+p$ connectors. Finally, we add Jordan arcs forming cycle of length $62d+$ 6 around the current arrangement. At the left and right end, this cycle is placed in such a way that every second arc of the cycle intersects a connector, in the correct order. The eight additional arcs of the cycle are used to connect the left and right side of the cycle, using four arcs on the top and four arcs on the bottom.
We denote the collection of all those arcs (including the pseudolines) as the enhanced pseudoline arrangement. The graph $G$ is given by the intersection graph of this arrangement. Note that $G$ could also be described purely combinatorially, and it can be constructed in polynomial time.
@@
This paragraph is dedicated to show that if $\mathcal{A}$ is stretchable then $G$ is realizable by unit segments. We assume that $\mathcal{A}$ is stretchable and show how to place each unit segment realizing $G$. We denote the segments representing the pseudolines, probes, connectors, and the cycle by important segments, probe segments, connector segments and cycle segments.
$\mathcal{A}$ being stretchable implies that there exists a combinatorially equivalent line arrangement $\mathcal{L}$. The arrangement $\mathcal{L}$ can be compressed (scaled down along the vertical axis) such that the slopes of all lines lie in some small interval $[-a,a]$, for say $a=1/20$. Additionally, we can move and scale $\mathcal{L}$ even more to ensure that all intersection points lie in the square $[-a,a]^{2}$. See Figure 9 for an illustration.
@@
For the next step we consider an important segment $s$ and construct its corresponding probe segments. We place all probe segments parallel to its important segment $s$, with sufficiently small distance to $s$ and to each other. The probe segments are placed as far towards the left as possible, while still reaching the intersections of important segments they need to reach. Since by construction of the scaled line arrangement $\mathcal{L}$ all intersections lie within $[-a,a]^{2}$ and the probes can only go until there, each probe segment and its corresponding important segment are almost collinear but shifted by roughly $0.5$ longitudinally See Figure 10 for an illustration of the placement of the probe segments.
Next we need to describe the placement of the connector segments. Note that on the right side, we only have important segments. We add all the connector segments in such a way that they lie on the same line as the important/probe segment they attach to. The connectors can overlap with the segments they attach to for a large part, since the first intersection of any probe and important segments only occurs in the square $[-a,a]^{2}$. This allows us to place our connectors such that all connectors on the left (right) side of the drawing end at the same left (right) $x$-coordinate.
Figure 10: Given $\mathcal{L}$, we can construct the probe segments and connector segments using unit segments. The connector segments are illustrated slightly tilted, for the purpose of readability. They should be parallel to the segment that they connect to.
@@
Finally, we can draw the cycle segments. For this, we simply make a sawtooth pattern on the left and on the right. See Figure 11 for an illustration. In our sawtooth pattern, every second cycle segment is horizontal, and all other cycle segments are parallel. The horizontal segments attach to the connector segments. As the important segments all have a very small slope, we can never run into a situation where two horizontal cycle segments would be too far away from each other to be connected. We connect the left and right sawtooth patterns to close the cycle, using our eight additional cycle segments.
@@
This paragraph is dedicated to show that $G$ being realizable using unit segments implies that $\mathcal{A}$ is stretchable. We thus assume there exists a realization $r$ of $G$ by unit segments. Similarly to the last paragraph, we denote the vertices representing the pseudolines, probes, connectors, and the cycle by important vertices $I$, probe vertices $P$, connector vertices $D$, and cycle vertices $C$.
@@
Note that $C$ forms an induced cycle in $G$ and $D$ are connectors of $C$ as in Definition 5 (thus motivating the name). By Lemma 6, $r(C)$ splits the plane into two cells, and $r(I\cup P)$ is contained completely in one of these two cells. Without loss of generality, we assume $r(I\cup P)$ is contained in the inner (bounded) cell, however all following arguments would also work with the outer cell. Thus, every segment in $r(D)$ intersects the cycle $r(C)$ from the inside, i.e., it intersects the interior boundary curve. We can thus apply Lemma 10, and get that $r(D)$ is ordered along the interior boundary curve of $C$ in the same way (up to cyclic shift and reflection) as it is in our enhanced pseudoline arrangement as described in the construction of $G$. Specifically, we know that the important segments and probe segments are ordered as in our enhanced pseudoline arrangement, see Figure 8.
We now claim that the arrangement of the important segments $r(I)$ is combinatorially equivalent to $\mathcal{A}$. Thus, extending these segments to lines yields a combinatorially equivalent line arrangement. To see this, we pick three pseudolines $\ell$, $a$, and $b$, such that $\ell$ intersects $a$ before $b$ when going from left to right in $\mathcal{A}$. We prove that in our realization $r$, the unit segment $r(\ell)$ must also intersect the segment $r(a)$ before $r(b)$. The orientation of the unit segments $r(\ell)$ is determined as follows: $r(\ell)$ intersects two connectors $d_{1},d_{2}$ that intersect $r(C)$ in two different unit segments. One of these unit segments $c$ (intersected by connector $d_{1}$) is two segments away from a unit segment that intersects a connector that intersects a probe of $\ell$. We then orient $r(\ell)$ in the direction such that the intersection with $d_{1}$ occurs before the intersection with $d_{2}$. Note that since $d_{1}$ and $d_{2}$ do not intersect, this order is uniquely defined. Furthermore, simply rotating or mirroring the representation does not change this order.
@@
Now, let us suppose for the purpose of a contradiction that $r(\ell)$ intersects $r(b)$ before $r(a)$. Let us consider the following curves:

### v3

@@ 2.1 Cell Lemma
We start by introducing the notion of a connector. Intuitively, a connector is just a set of vertices that separates an induced cycle from the (connected) rest of the graph.
@@
Let $G=(V,E)$ be a graph, and let $C\subseteq V$ be a set of vertices forming an induced cycle such that $G[V\setminus C]$ is connected. Let $D\subseteq V\setminus C$ be the neighborhood of $C$. If all of the following properties hold, we call $D$ the set of connectors of $C$.
@@
This definition is illustrated in Figure˜7.
@@
Let $C$ be an induced cycle of $G=(V,E)$ with connectors $D$, and let $r(V)$ be a geometric representation of $G$ by Jordan arcs. Then $r(V\setminus(C\cup D))$ lies completely in one cell of the arrangement $r(C)$.
@@
This follows from the fact that $G[V\setminus(C\cup D)]$ is connected and is not adjacent to the induced cycle $C$. Therefore, $r(V\setminus(C\cup D))$ is a connected subset of the plane that does not intersect $r(C)$, and so completely lies in one of its cells. ∎
Finally, note that if in the previous lemma $|D|\geq 5$, then the precondition of Lemma˜3 is also satisfied.
Figure 7: Left: The yellow dots form an induced cycle. The green vertices are connectors of this cycle and the remaining white squares form the rest of the graph. Right: The yellow dots form an induced cycle. The blue dots are intersectors of this cycle. The cyclic order $o_{G}$ as used in Lemma˜9 is $d_{2},d_{2},d_{1},d_{4},d_{3},d_{1}$, when starting at the top left of the cycle and going clockwise.
@@ 2.2 Order Lemma
In this section we show that certain vertices neighboring an induced cycle also need to intersect this cycle in the same order in any geometric representation. We start with a definition of the setup.
@@
Let $G=(V,E)$ be a graph, and let $C\subseteq V$ be a set of vertices forming an induced cycle. Let $D\subseteq V\setminus C$ be a set of vertices with the following properties.
@@
  • If $d\in D$ has two neighbors in $C$, then these are non-adjacent.
@@
Then we call $D$ a set of intersectors of $C$.
This definition is illustrated in Figure˜7. We now define two cyclic orders of intersections of intersectors with the cycle. The first order lives in the realm of the graph, while the second one is concerned with a concrete realization. The goal of this section is to prove that these orders are the same.
@@
Let $G$ be a graph with $C$ an induced cycle on at least four vertices, and let $D$ be a set of intersectors of $C$. When we travel along the cycle $C$ for one full rotation, we write down the pairs $(c,d)\in C\times D$ such that $\{c,d\}\in E(G)$. This sequence of pairs defines the graph order of the intersectors up to a cyclic shift and reflection.
@@
Let $G$ be a graph with $C$ an induced cycle on at least four vertices, and let $D$ be a set of intersectors of $C$. Let $r$ be a realization of $G$ by unit segments, and let $b$ be the interior boundary curve of $r(C)$. More generally, if $r(C)$ is a generalization with Jordan curves, let $b$ be the boundary curve of any cell $F$ such that $\operatorname{trace}(F)$ has at least 5 distinct symbols. When we travel along $b$ for one full rotation, we write down the pairs $(c,d)\in C\times D$ such that $r(d)$ intersects $b$ where $b$ is tracing the boundary of the cell $F$ in $r(c)$. From consecutive copies of the same pair we only keep one. This sequence of pairs defines the geometric order of the intersectors up to a cyclic shift and reflection.
The graph order is illustrated in Figure˜7, while the geometric order is illustrated in Figure˜8.
@@
Let $r$ be a realization of a graph $G$ with an induced cycle $C$ and intersectors $D$. If for every intersection of some $r(d)$ for $d\in D$ with some $r(c)$ for $c\in C$ we have that $r(d)$ also intersects the boundary curve $b$ close to that intersection, then the geometric order and the graph order of the intersectors coincide.
Figure 8: A realization $r$ of an induced cycle $C$ (yellow Jordan arcs) and intersectors $\{d_{1},d_{2}\}$ of $C$ (green Jordan arcs). The cyclic order $o_{r}$ along the interior boundary curve as used in Lemma˜9 is $(c_{5},d_{2}),(c_{7},d_{1}),(c_{2},d_{1})$, when starting at the bottom right and going clockwise (see black arrow).
@@
We first claim that every pair $(c,d)$ occurs in both orders exactly once if $\{c,d\}\in E(G)$, and zero times otherwise: In the graph order this holds by definition. Furthermore, the assumption of this lemma guarantees that any pair $(c,d)$ in the graph order of intersectors also occurs in the geometric order of intersectors. We thus need to show only that in the geometric order no pair occurs multiple times. A pair could occur more than once only if there is another pair $(c^{\prime},d^{\prime})$ between these occurrences. However, since $c$ is crossed only by $d$, and no neighbor of $c$ in $C$ is crossed by any intersector, Lemma˜3 guarantees that no such pair $(c^{\prime},d^{\prime})$ can occur between the two occurrences of $(c,d)$.
Next, we show that the pairs are also ordered the same way. Since for any two pairs $(c,d)$ and $(c^{\prime},d^{\prime})$ in the order the two vertices $c,c^{\prime}$ are distinct and non-adjacent, Lemma˜3 guarantees that the geometric order of the intersectors must respect the ordering of the $c_{i}$ along the cycle. Similarly, the graph order of intersectors must respect this ordering by definition. Thus, the two orders are the same. ∎
Note that the definition of the geometric order and Lemma˜9 also work for the exterior boundary curve instead of the interior one.
@@ 3 $\exists\mathbb{R}$-Membership
In this section, we show the $\exists\mathbb{R}$-membership parts of Theorems˜1 and 2.
@@
In order to present the naive approach, we restrict our attention to only unit segments. However a similar technique also easily works for polylines. Let $G=(V,E)$ be a graph. For each vertex $v\in V$, we use four variables $\left\llbracket v1\right\rrbracket,\left\llbracket v2\right\rrbracket,\left\llbracket v3\right\rrbracket,\left\llbracket v4\right\rrbracket$ to describe the coordinates of the endpoints of a unit segment realizing $v$. We can construct ETR-formulas unit and intersection that test whether a segment has unit length and whether two segments intersect, respectively. The formula $\varphi$ for Unit Segment Recognition consists of the three parts:
@@
The unit formula can be constructed using the formula for the Euclidean norm. To construct the intersection formula it is possible to use the orientation test: The orientation test formula checks whether a given ordered triple of points is oriented clockwise, or counter-clockwise. Using multiple orientation tests on the endpoints of two segments, one can determine whether the segments intersect. The orientation test itself can be constructed using a standard determinant test. This finishes the description of the formula $\varphi$ and establishes $\exists\mathbb{R}$-membership of Unit Segment Recognition.
Although all of these formulas are straightforward to describe, things get a bit lengthy (especially in the case of polylines) and we do hide some details about the precise polynomials. We therefore also describe the second approach using real witnesses and verification algorithms. To use this approach we need to first introduce a different characterization of the complexity class $\exists\mathbb{R}$. Namely, an algorithmic problem is in $\exists\mathbb{R}$ if and only if we can provide a real verification algorithm [19]. A real verification algorithm $A$ for a problem $P$ takes as input an instance $I$ of $P$ and a polynomial-size real-valued witness $w$. AA must have the following properties: In case that $I$ is a yes-instance, there exists a $w$ such that $A(I,w)$ returns yes. In case that $I$ is a no-instance, $A(I,w)$ will return no, for all possible $w$. Note that this is reminiscent of the definition of NP using a verifier algorithm. There are two subtle differences: The first one is that $w$ is allowed to contain real numbers and discrete values, not only bits. The second difference is that $A$ runs on the real RAM instead of a Turing machine. This is required since a Turing machine is not capable of dealing with real numbers. It is important to note that $I$ itself does not contain any real numbers. We refer to Erickson, van der Hoog, and Miltzow [19] for a detailed definition of the real RAM.
Given this alternative characterization of $\exists\mathbb{R}$, it is now very easy to establish $\exists\mathbb{R}$-membership of Unit Segment Recognition and $k$-PolyLine Recognition: We merely need to describe the witness and the verification algorithm. The witness is a description of the coordinates of the unit segments (or polylines, respectively) realizing the given graph. The verification algorithm checks that the realizations of vertices do or do not intersect one another, and in the case of unit segments, also checks that all segments have the correct length. Note that we sweep many details of the algorithm under the carpet. However, algorithms are much more versatile than formulas and it is a well-established fact that algorithms are capable of all types of elementary operations needed to perform this verification.
@@ 4 $\exists\mathbb{R}$-Hardness
In this section we prove $\exists\mathbb{R}$-hardness, first for unit segments, then for polylines. The reduction for polylines builds upon the reduction for unit segments, and we will only highlight the differences.
@@ 4.1 $\exists\mathbb{R}$-Hardness for Unit Segments
We show $\exists\mathbb{R}$-hardness of Unit Segment Recognition by a reduction from Stretchability.
@@
We present the reduction in three steps. First, we show how we construct a graph $G$ from a pseudoline arrangement $\mathcal{A}$. Next we show completeness, i.e., we show that if $\mathcal{A}$ is stretchable then $G$ can be represented using unit segments. At last, we show soundness, i.e., we show that if $G$ can be represented using unit segments then $\mathcal{A}$ is stretchable. For this part we will use the lemmas from Section˜2.
@@
Given a pseudoline arrangement $\mathcal{A}$ of $n$ pseudolines, we construct a graph $G$ by enhancing the arrangement $\mathcal{A}$ with additional Jordan arcs. Then we define $G$ to be the intersection graph of the pseudolines (which are also Jordan arcs) and all our added arcs. See Figure˜9 for an illustration.
First, we add so called probes to our pseudolines. A probe of depth $k$ of pseudoline $\ell$ is a Jordan arc that starts on the left vertical line, and follows $\ell$ closely through the arrangement until it has intersected $k$ other pseudolines (and all probes of other pseudolines that reach this intersection). For each pseudoline $\ell$, we add 2(n1)2(n-1) probes: Above and below $\ell$, we add one probe each for each depth $1\leq k\leq n-1$. This gives a total of p:=2n(n1)p:=2n(n-1) probes. The probes are sorted according to their depth, with the probes of smallest depth situated closest to $\ell$. Note that so far we have added $n+p$ Jordan arcs. We now create connectors for each probe and pseudoline. A connector is a short Jordan arc added to the left and/or right end of another arc. For probes, we only add connectors at the left end, while pseudolines get connectors at both ends. Note that we thus add $d:=2n+p$ connectors. Finally, we add Jordan arcs forming a cycle of length $2d+6$ around the current arrangement. On the left and right side, this cycle is placed in such a way that every second arc of the cycle intersects a connector, in the correct order. The eight remaining arcs of the cycle connect the left and right side of the cycle, using four arcs on the top and four arcs on the bottom.
We call the collection of all these arcs (including the pseudolines) the enhanced arrangement. The graph $G$ is the intersection graph of this arrangement. Note that $G$ could also be described purely combinatorially, and it can be constructed in polynomial time.
@@
In this paragraph, we show that if $\mathcal{A}$ is stretchable then $G$ is realizable by unit segments. We assume that $\mathcal{A}$ is stretchable and show how to place each unit segment realizing $G$. We call the segments representing the pseudolines, probes, connectors, and the cycle important segments, probe segments, connector segments, and cycle segments, respectively.
As the pseudoline arrangement $\mathcal{A}$ is stretchable, there exists a combinatorially equivalent line arrangement $\mathcal{L}$. The arrangement $\mathcal{L}$ can be compressed (scaled down along the vertical axis) such that the slopes of all lines lie in some small interval [a,a][-a,a] for, say, $a=1/20$. Additionally, we can move and scale $\mathcal{L}$ even more to ensure that all intersection points lie in the square $a,a]^{2}$. See Figure˜10 for an illustration.
@@
For the next step we consider an important segment $s$ and construct its corresponding probe segments. We place all probe segments parallel to its important segment $s$, with sufficiently small distance to $s$ and to each other. The probe segments are placed as far towards the left as possible, while still reaching the intersections of important segments they need to reach. Since by construction of the scaled line arrangement $\mathcal{L}$, all intersections lie within $a,a]^{2}$ and the probes can go only until there, each probe segment and its corresponding important segment are almost collinear but shifted by roughly $0.5$ longitudinally. See Figure˜11 for an illustration of the placement of the probe segments.
Next we need to describe the placement of the connector segments. Note that on the right side, we only have important segments. We add all the connector segments in such a way that they lie on the same line as the important/probe segment they attach to. The connectors can overlap with the segments they attach to for a large part since the first intersection of any probe and important segments only occurs in the square $a,a]^{2}$. This allows us to place our connectors such that all connectors on the left (right) side of the drawing end at the same left (right) $x$-coordinate.
Figure 11: Given the line arrangement $\mathcal{L}$, we can construct the probe segments and connector segments using unit segments. The connector segments are illustrated slightly tilted, for the purpose of readability. They should be parallel to the segment that they connect to.
@@
Finally, we can draw the cycle segments. For this, we simply make a sawtooth pattern on the left and on the right. See Figure˜12 for an illustration. In our sawtooth pattern, every second cycle segment is horizontal, and all other cycle segments have the same, slightly positive slope. The horizontal segments attach to the connector segments. As the important segments all have a very small slope, we can never run into a situation where two horizontal cycle segments would be too far away from each other to be connected. We connect the left and right sawtooth patterns to close the cycle, using our eight additional cycle segments. (For the following reason, eight segments are sufficient for this task: Since the left and right sawtooth patterns are connected by a path that consists of a left connector segment, an important segment, and a right connector segment, their distance is at most $3$, allowing us to form both the top and the bottom boundary of the cycle with at most four cycle segments each.)
@@
In this paragraph, we show that $G$ being realizable using unit segments implies that $\mathcal{A}$ is stretchable. We thus assume that there exists a realization $r$ of $G$ by unit segments. Similarly to the last paragraph, we denote the sets of vertices representing the pseudolines, probes, connectors, and the cycle by important vertices $I$, probe vertices $P$, connector vertices $D$, and cycle vertices $C$.
@@
Note that $C$ forms an induced cycle in $G$ and $D$ is a set of connectors of $C$ as in Definition˜4 (thus motivating the name). By Lemma˜5, $r(C)$ splits the plane into two cells, and $r(I\cup P)$ is contained completely in one of these two cells. Without loss of generality, we assume $r(I\cup P)$ is contained in the inner (bounded) cell, however all following arguments would also work with regards to the outer cell. Thus, every segment in $r(D)$ intersects the cycle $r(C)$ from the inside, i.e., it intersects the interior boundary curve. We can thus apply Lemma˜9, and get that $r(D)$ is ordered along the interior boundary curve of $C$ in the same way (up to cyclic shift and reflection) as it is in the enhanced arrangement from which we defined $G$. Specifically, we know that the important segments and probe segments are ordered as in our enhanced arrangement; see Figure˜9.
We now claim that the arrangement of the important segments $r(I)$ is combinatorially equivalent to $\mathcal{A}$. In order to show this claim, our proof strategy is to define two orientations: One orientation $O_{\mathcal{A}}$ of the pseudolines in the pseudoline arrangement $\mathcal{A}$, and one orientation $O_{r}$ of the important segments $r(I)$ in the realization $r$. We show that the order in which pseudolines intersect other pseudolines with respect to $O_{\mathcal{A}}$ is equal to the order in which important segments are intersected by other important segments with respect to $O_{r}$. If we can show this claim, we are done, since the order of important segments and unit segments along the interior boundary in $r(I)$ is the same as in our enhanced arrangement. Therefore, (by extending the important unit segments $I$ to infinite lines) we obtain that $\mathcal{A}$ is stretchable.
It remains to show the claim. The orientations $O_{\mathcal{A}}$ and $O_{r}$ are defined as follows: In the orientation $O_{\mathcal{A}}$, every pseudoline is simply oriented from left to right (where left and right correspond to the directions in Figure˜9). The orientation $O_{r}$ is defined as follows: Consider some important vertex $v\in I$, and its corresponding important segment $r(v)$. Note that $v$ has two neighbors $d_{1},d_{2}$ in $G$ such that $d_{1},d_{2}\in D$ are connector vertices. We can assume that $d_{1}$ is a connector segment corresponding to a connector on the left (of Figure˜9) and $d_{2}$ is a connector segment corresponding to a connector on the right (of Figure˜9). Since $v,d_{1}$ are neighbors in $G$, the unit segment $r(v)$ has a crossing $c_{1}$ with $r(d_{1})$. Likewise it has a crossing $c_{2}$ with $r(d_{1})$. In the orientation $O_{r}$ we orient the segment $r(v)$ such that $c_{1}$ comes before $c_{2}$. (Note that if the unit segments $r(d_{i}),r(v)$ are not parallel for $i=1,2$ then the crossings $c_{1},c_{2}$ are unique and $c_{1}\neq c_{2}$, and so $O_{r}$ is well-defined. In the case that $r(d_{i}),r(v)$ are parallel it is still well-defined since $r(d_{1}),r(d_{2})$ do not intersect.)
Now we are ready to show the claim. We pick three pseudolines $\ell$, $a$, and $b$ such that $\ell$ intersects $a$ before $b$ in the orientation $O_{\mathcal{A}}$. We prove that in the orientation $O_{r}$, the unit segment $r(\ell)$ must also intersect the segment $r(a)$ before $r(b)$. (In order to simplify the notation, we use the same symbol for a pseudoline, as well as its corresponding vertex in $G$).
@@
Now, suppose for the purpose of a contradiction that $r(\ell)$ intersects $r(b)$ before $r(a)$. Consider the following curves: